OpenStudy (anonymous):

Solve (4x^2-4x+1) / (8x^3-1)

OpenStudy (anonymous):

what does "solve" mean in this case?

OpenStudy (anonymous):

Divide

OpenStudy (anonymous):

note that 4x^2-4x+1 = (2x-1)^2 and 8x^3 -1 = ((2x)^3 - 1^3)

OpenStudy (anonymous):

in other words factor the numerator and denominator and cancel. numerators i s \[(2x-1)(2x-1)\] denominator os \[(2x-1)(4x^2+2x+1)\]

OpenStudy (anonymous):

giving \[\frac{2x-+1}{4x^2+2x+1}\]